### v1

@@ Abstract
Message passing algorithms have been tailored for compressive imaging applications by plugging in different types of off-the-shelf image denoisers. These off-the-shelf denoisers mostly rely on some generic or hand-crafted priors for denoising. Due to their insufficient accuracy in capturing the true image prior, these methods often fail to produce satisfactory results, especially in largely underdetermined scenarios. On the other hand, score-based generative modeling offers a promising way to accurately characterize the sophisticated image distribution. In this paper, by exploiting the close relation between score-based modeling and empirical Bayes-optimal denoising, we devise a message passing framework that integrates a score-based minimum mean squared error (MMSE) denoiser for compressive image recovery. This framework is firmly rooted in Bayesian formalism, in which state evolution (SE) equations accurately predict its asymptotic performance. Experiments on the FFHQ dataset demonstrate that our method strikes a significantly better performance-complexity tradeoff than conventional message passing, regularized linear regression, and score-based posterior sampling baselines. Remarkably, our method typically requires less than $20$ neural function evaluations (NFEs) to converge.
Keywords: Message passing, score-based generative models, plug-and-play priors, linear inverse problems
1 Introduction
@@
Inspired by belief propagation (BP) in graphical models, message passing algorithms stand as a powerful tool for Bayesian compressive sensing (CS), enabling the incorporation of a variety of sparsity-inducing priors. To name a few, approximate message passing (AMP) [9] utilizes a quadratic approximation of max-sum loopy BP to efficiently implement the least absolute shrinkage and selection operator (LASSO). The LASSO solution is equivalent to the maximum a posteriori (MAP) estimator assuming a Laplacian prior. Turbo-CS [21], also known as vector approximate message passing (VAMP) [25], employs a Bernoulli-Gaussian distribution for modeling the sparsity pattern. Please refer to [37] for a comprehensive review of other variants of message passing algorithms.
@@
Unfortunately, sparsity-based approaches often exhibit poor performance in many imaging applications. The reason is that natural images do not have an exact sparse representation in any known transform domain (e.g., discrete cosine transform (DCT), wavelet, curvelet). To address this limitation, off-the-shelf image denoisers have been plugged into message passing algorithms for performance enhancement [24, 27, 34]. Popular choices include the total variation (TV) denoiser, the SURE-LET denoiser [3], the BM3D denoiser [7], etc. These off-the-shelf denoisers mostly rely on non-parametric models for denoising, which do not require the hypothesis of a statistical model for the noiseless image. Consequently, the above-mentioned plug-and-play methods can result in sub-optimal performance, as they fail to fully exploit the strong prior inherent in image data. Some recent works [23, 32] also incorporate either a denoising CNN (DnCNN) or an unrolled network in place of the off-the-shelf denoiser.
@@
On the other hand, score-based generative models [29, 15, 30] have achieved unparalleled success in generating high-quality samples across various domains, including images, videos, and molecular conformations. This evidences their remarkable capabilities in capturing the intricacies of high-dimensional data distribution. During training, score-based generative models perturb the data with random Gaussian noise of different magnitudes. They learn the data prior by matching the gradient of the log density of the perturbed data, known as score function. The score function is then applied at each and every reverse diffusion step for sample generation.
The score-based generative models can also be leveraged for solving inverse problems without task-specific training [17, 18, 5, 6, 28, 36, 10]. These methods obtain a prior-related term (i.e., the score function) from the score-based generative network and a likelihood term from the degradation model. The two terms are combined to form the posterior term via Bayes’ theorem, enabling solving inverse problems stochastically by sampling from the posterior distribution. However, these methods typically require traversing the entire reverse diffusion process, necessitating hundreds or even thousands of neural function evaluations (NFEs), which makes them computationally expensive. Moreover, sampling from the exact posterior distribution is indeed intractable due to the time dependence of the likelihood. This challenge has led to various approximations used in posterior sampling, typically achieved through projection at each time step to enforce data consistency [5], or a reweighted combination of the prior score and the estimated likelihood function [6]. Please refer to a recent survey paper [8] for different categories of approximations. These approximations often introduce additional hyperparameters that are sensitive and require careful tuning. Furthermore, the approximation errors accumulate through successive iterations, making it challenging to assess their performance theoretically.
@@
In this paper, we devise a score-based message passing algorithm for compressive imaging within the Bayesian framework. This approach inherits the fast convergence speed of message passing and, meanwhile, fully leverages the image prior through the integration of score-based networks. We exploit the fact that, at each iteration, message passing algorithms construct a pseudo AWGN observation of the ground-truth image. We repurpose the score function for denoising the AWGN observation through Tweedie’s formula [11], achieving per-iteration minimum mean squared error (MMSE) performance with a single forward pass through the score-based network. This significantly differs from the typical usage of the score function for reverse diffusion sampling. We further relate the posterior variance calculation to the second-order generalization of Tweedie’s formula, which depends on the Hessian of the log density (i.e., the second-order score function) [22]. We propose matching the second-order score function to obtain the posterior variance, facilitating message updates in a turbo-type message passing algorithm. This approach leads to the name score-based turbo message passing (STMP). We present the state evolution (SE) analysis to track the asymptotic MSE performance of STMP. Experimental results on the FFHQ dataset [16] demonstrate that STMP strikes a much better tradeoff between performance and complexity than existing conventional message passing, regularized linear regression, and score-based posterior sampling schemes. Notably, STMP typically requires less than $20$ NFEs to converge.
2 Turbo-CS Framework
In this section, we review the Turbo-CS framework [21, 25] based on an expectation-propagation (EP) type message-passing derivation on the factor graph.
Figure 1: Factor graph for Turbo-CS derivation. The circles represent variable nodes and the squares represent factor nodes.
The posterior distribution can be factorized as
$p(\mathbf{x}|\mathbf{y})\propto p(\mathbf{x})p(\mathbf{y}|\mathbf{x}).$ (2)
By letting $\mathbf{x}_{\sf A}=\mathbf{x}_{\sf B}=\mathbf{x}$, we obtain an equivalent factorization of (2):
$p(\mathbf{x}_{\sf A},\mathbf{x}_{\sf B}|\mathbf{y})\propto p(\mathbf{x}_{\sf B})\delta(\mathbf{x}_{\sf A}-\mathbf{x}_{\sf B})p(\mathbf{y}|\mathbf{x}_{\sf A}),$ (3)
where $\delta(\cdot)$ is the Dirac delta function. The corresponding factor graph representation is shown in Fig. 1.
We start by initializing the message from the factor node $\delta(\mathbf{x}_{\sf A}-\mathbf{x}_{\sf B})$ to the variable node $\mathbf{x}_{\sf A}$ as $\mu_{\delta\rightarrow\sf{A}}(\mathbf{x}_{\sf A})=\mathcal{N}\big{(}\mathbf{x}_{\sf A};\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}}\mathbf{I}\big{)}$. According to the EP approximation, the belief on $\mathbf{x}_{\sf A}$ (i.e., the marginal posterior density estimate of $\mathbf{x}_{\sf A}$) is given by $b(\mathbf{x}_{\sf A})=\mathcal{N}\big{(}\mathbf{x}_{\sf A};\mathbf{x}_{\sf A}^{\textit{post}},\mathbf{V}_{\sf A}^{\textit{post}}\big{)}=\mathrm{proj}\left[p(\mathbf{y}|\mathbf{x}_{\sf A})\mu_{\delta\rightarrow\sf{A}}(\mathbf{x}_{\sf A})\right]$, where $\mathrm{proj}[p(\cdot)]$ denotes the projection of $p(\cdot)$ to a Gaussian probability density function (PDF) with matched first- and second-order moments. By noting that the product of two Gaussian messages is proportional to a Gaussian PDF, the projection operator in $b(\mathbf{x}_{\sf A})$ can thus be removed. We further approximate $\mathbf{V}_{\sf A}^{\textit{post}}$ using a shared scalar variance [21, 25], resulting in the belief of the form $b(\mathbf{x}_{\sf A})\approx\mathcal{N}\big{(}\mathbf{x}_{\sf A};\mathbf{x}_{\sf A}^{\textit{post}},v_{\sf A}^{\textit{post}}\mathbf{I}\big{)}$, where
@@ 2 Turbo-CS Framework
Eqns. (4) and (5) are respectively the posterior mean and variance from the linear minimum mean-square error (LMMSE) estimator of $\mathbf{x}_{\sf A}$, given the prior mean $\mathbf{x}_{\sf A}^{\textit{pri}}$ and variance $v_{\sf A}^{\textit{pri}}$. The message from the variable node $\mathbf{x}_{\sf A}$ to the factor node $\delta(\mathbf{x}_{\sf A}-\mathbf{x}_{\sf B})$, also known as the extrinsic message, is calculated as $\mu_{\sf{A}\rightarrow\delta}(\mathbf{x}_{\sf A})=\mathcal{N}\big{(}\mathbf{x}_{\sf A};\mathbf{x}_{\sf A}^{\textit{ext}},v_{\sf A}^{\textit{ext}}\mathbf{I}\big{)}=\mathcal{N}\big{(}\mathbf{x}_{\sf A};\mathbf{x}_{\sf A}^{\textit{post}},v_{\sf A}^{\textit{post}}\mathbf{I}\big{)}/\mathcal{N}\big{(}\mathbf{x}_{\sf A};\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}}\mathbf{I}\big{)}$, yielding Lines 7 and 8 of Algorithm 1. The extrinsic message decorrelates the input and output estimation errors, i.e., $\mathbb{E}\big{[}(\mathbf{x}_{\sf A}-\mathbf{x}_{\sf A}^{\textit{pri}})(\mathbf{x}_{\sf A}-\mathbf{x}_{\sf A}^{\textit{ext}})^{\sf T}\big{]}=\mathbf{0}$ [20].
Algorithm 1 Turbo-CS Framework
2:  Output: $\mathbf{x}_{\sf B}^{\textit{post}}$
3:  repeat
4:     % LMMSE estimator
10:     % MMSE denoiser
11:     $\mathbf{x}_{\sf B}^{\textit{post}}=\mathbb{E}\big{[}\mathbf{x}_{\sf B}\big{|}\mathbf{x}_{\sf B}^{\textit{pri}}\big{]}$
12:     $v_{\sf B}^{\textit{post}}=\frac{1}{N}\mathrm{tr}\big{(}\mathrm{Cov}\big{[}\mathbf{x}_{\sf B}\big{|}\mathbf{x}_{\sf B}^{\textit{pri}}\big{]}\big{)}$
16:  until the stopping criterion is met
The message $\mu_{\sf{A}\rightarrow\delta}(\mathbf{x}_{\sf A})=\mathcal{N}\big{(}\mathbf{x}_{\sf A};\mathbf{x}_{\sf A}^{\textit{ext}},v_{\sf A}^{\textit{ext}}\mathbf{I}\big{)}$ flows rightward through the factor node $\delta(\mathbf{x}_{\sf A}-\mathbf{x}_{\sf B})$ unchanged, manifesting as $\mu_{\delta\rightarrow\sf{B}}(\mathbf{x}_{\sf B})=\mathcal{N}\big{(}\mathbf{x}_{\sf B};\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}}\mathbf{I}\big{)}$ on the other side. This leads to Line 9 of Algorithm 1. Then, based on the EP approximation, the belief on $\mathbf{x}_{\sf B}$ is calculated as $b(\mathbf{x}_{\sf B})=\mathcal{N}\big{(}\mathbf{x}_{\sf B};\mathbf{x}_{\sf B}^{\textit{post}},\mathbf{V}_{\sf B}^{\textit{post}}\big{)}=\mathrm{proj}\left[p(\mathbf{x}_{\sf B})\mu_{\delta\rightarrow\sf{B}}(\mathbf{x}_{\sf B})\right]$, where
$\mathbf{x}_{\sf B}^{\textit{post}}$ $=\frac{1}{Z}\int_{\mathbf{x}_{\sf B}}\mathbf{x}_{\sf B}p(\mathbf{x}_{\sf B})\mathcal{N}\big{(}\mathbf{x}_{\sf B};\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}}\mathbf{I}\big{)},$ (6)
$\mathbf{V}_{\sf B}^{\textit{post}}$ $=\frac{1}{Z}\int_{\mathbf{x}_{\sf B}}\big{(}\mathbf{x}_{\sf B}-\mathbf{x}_{\sf B}^{\textit{pri}}\big{)}\big{(}\mathbf{x}_{\sf B}-\mathbf{x}_{\sf B}^{\textit{pri}}\big{)}^{\sf T}p(\mathbf{x}_{\sf B})\mathcal{N}\big{(}\mathbf{x}_{\sf B};\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}}\mathbf{I}\big{)},$ (7)
with $Z=\int_{\mathbf{x}_{\sf B}}p(\mathbf{x}_{\sf B})\mathcal{N}\big{(}\mathbf{x}_{\sf B};\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}}\mathbf{I}\big{)}$ being the normalization constant. The state evolution (SE) analysis [13] has shown that the error vector $\mathbf{x}_{\sf B}^{\textit{pri}}-\mathbf{x}_{\sf B}$ is asymptotically i.i.d. Gaussian as $M,N\rightarrow\infty$ with a fixed ratio. Based on this, the Turbo-CS algorithm models $\mathbf{x}_{\sf B}^{\textit{pri}}$ as an AWGN observation of the true signal $\mathbf{x}_{\sf B}$, i.e.,
$\mathbf{x}_{\sf B}^{\textit{pri}}=\mathbf{x}_{\sf B}+\mathbf{w},~{}~{}\mathbf{w}\sim\mathcal{N}\big{(}\mathbf{0},v_{\sf B}^{\textit{pri}}\mathbf{I}\big{)}.$ (8)
From (8), we have $p\big{(}\mathbf{x}_{\sf B}^{\textit{pri}}|\mathbf{x}_{\sf B}\big{)}=\mathcal{N}\big{(}\mathbf{x}_{\sf B}^{\textit{pri}};\mathbf{x}_{\sf B},v_{\sf B}^{\textit{pri}}\mathbf{I}\big{)}$. By noting that $\mathcal{N}\big{(}\mathbf{x}_{\sf B};\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}}\mathbf{I}\big{)}$ in (6) and (7) takes the same form as $p\big{(}\mathbf{x}_{\sf B}^{\textit{pri}}|\mathbf{x}_{\sf B}\big{)}$, the belief calculation on $\mathbf{x}_{\sf B}$ can be interpreted as the MMSE denoising under the AWGN observation model (8):
$\mathbf{x}_{\sf B}^{\textit{post}}$ $=\mathbb{E}\big{[}\mathbf{x}_{\sf B}\big{|}\mathbf{x}_{\sf B}^{\textit{pri}}\big{]},$ (9)
$\mathbf{V}_{\sf B}^{\textit{post}}$ $=\mathrm{Cov}\big{[}\mathbf{x}_{\sf B}\big{|}\mathbf{x}_{\sf B}^{\textit{pri}}\big{]}.$ (10)
Similarly, we approximate $\mathbf{V}_{\sf B}^{\textit{post}}$ by $v_{\sf B}^{\textit{post}}\mathbf{I}$ with $v_{\sf B}^{\textit{post}}=\frac{1}{N}\mathrm{tr}\big{(}\mathrm{Cov}\big{[}\mathbf{x}_{\sf B}\big{|}\mathbf{x}_{\sf B}^{\textit{pri}}\big{]}\big{)}$ to simplify calculation. The extrinsic message of $\mathbf{x}_{\sf B}$ is denoted as $\mu_{{\sf B}\rightarrow\delta}(\mathbf{x}_{\sf B})=\mathcal{N}\big{(}\mathbf{x}_{\sf B};\mathbf{x}_{\sf B}^{\textit{ext}},v_{\sf B}^{\textit{ext}}\mathbf{I}\big{)}$, where the variance and mean are respectively given in Lines 13 and 14 of Algorithm 1. Finally, the extrinsic message of $\mathbf{x}_{\sf B}$ flows leftward through the factor node $\delta(\mathbf{x}_{\sf A}-\mathbf{x}_{\sf B})$ unchanged, serving as the prior message of $\mathbf{x}_{\sf B}$ in the next iteration, i.e., $\mathbf{x}_{\sf A}^{\textit{pri}}=\mathbf{x}_{\sf B}^{\textit{ext}},v_{\sf A}^{\textit{pri}}=v_{\sf B}^{\textit{ext}}$.
Figure 2: Modularized representation of the Turbo-CS algorithm.
As illustrated in Fig. 2, the above message passing principle admits a modularized interpretation that resembles the turbo decoder [1], hence the name Turbo-CS. In particular, the Turbo-CS framework consists of two modules. Module A is a simplified LMMSE estimator of $\mathbf{x}$ based on the observation $\mathbf{y}$, in which the messages from module B serve as the prior mean and variance. With the equivalent AWGN observation model (8), Module B can be interpreted as an MMSE denoiser incorporating the prior distribution of $\mathbf{x}$ and the messages from module A. The two modules are executed iteratively to refine the estimate of $\mathbf{x}$.
@@
The main challenge in implementing the Turbo-CS algorithm arises from the MMSE denoising part in module B, primarily due to i) the lack of a tractable prior distribution and ii) the high-dimensional integrals involved in calculating the posterior mean and variance. To sidestep these difficulties, various off-the-shelf image denoisers (e.g., TV, SURE-LET [3], BM3D [7]) are plugged into message passing algorithms in place of the principled MMSE denoiser. These plug-and-play denoisers typically rely on generic or hand-crafted priors, which could lead to substantial imperfection in compressive imaging applications due to oversimplification and misspecification of the prior distribution. Motivated by the above, we propose a score-based turbo message passing (STMP) algorithm that exploits the score-based models as an accurate and expressive image prior, facilitating compressive image recovery from a purely Bayesian perspective.
3 Score-Based MMSE Denoising
Theorem 1 (Tweedie’s formula and its second-order generalization).

### v2

@@ Abstract
Message passing algorithms have been tailored for compressive imaging applications by plugging in different types of off-the-shelf image denoisers. These off-the-shelf denoisers mostly rely on some generic or hand-crafted priors for denoising. Due to their insufficient accuracy in capturing the true image prior, these methods often fail to produce satisfactory results, especially in highly underdetermined scenarios. On the other hand, score-based generative modeling offers a promising way to accurately characterize the sophisticated image distribution. In this paper, by exploiting the close relation between score-based modeling and empirical Bayes-optimal denoising, we devise a message passing framework that integrates a score-based minimum mean squared error (MMSE) denoiser for compressive image recovery. Experiments on the FFHQ dataset demonstrate that our method strikes a significantly better performance-complexity tradeoff than conventional message passing, regularized linear regression, and score-based posterior sampling baselines. Remarkably, our method typically converges in fewer than 20 neural function evaluations (NFEs).
I Introduction
@@
Message passing algorithms [1, 2, 3] stand as a powerful tool for compressive image recovery. For example, denoising-based approximate message passing (D-AMP) [1] and denoising-based turbo compressive sensing (D-Turbo-CS) [3] are able to incorporate a variety of off-the-shelf image denoisers. Popular choices include the total variation (TV) denoiser, the SURE-LET denoiser [4], the BM3D denoiser [5], etc. These off-the-shelf denoisers mostly rely on non-parametric models for denoising, which do not require the hypothesis of a statistical model for the noiseless image. Consequently, the above-mentioned plug-and-play methods can result in sub-optimal performance, as they fail to fully exploit the strong prior inherent in image data. Some recent works [6, 7] incorporate either a denoising CNN (DnCNN) or an unrolled network in place of the off-the-shelf denoiser.
@@
On the other hand, score-based generative models [8] have achieved unparalleled success in generating high-quality samples across various domains. This evidences their remarkable capabilities in capturing the intricacies of high-dimensional data distribution. During training, score-based generative models perturb the data with random Gaussian noise of different magnitudes. They learn the data prior by matching the gradient of the log density of the perturbed data, known as score function. The score function is then applied at each and every reverse diffusion step for sample generation. Score-based generative models can also be leveraged for solving inverse problems without task-specific training [9, 10, 11]. These methods obtain a prior-related term (i.e., the score function) from the score-based generative network and a likelihood term from the degradation model. The two terms are combined to form the posterior term via Bayes’ theorem, enabling solving inverse problems stochastically by sampling from the posterior distribution. However, these methods typically require traversing the entire reverse diffusion process, necessitating hundreds or even thousands of neural function evaluations (NFEs), which makes them computationally expensive.
@@
In this paper, we devise a score-based message passing algorithm for compressive imaging. This approach inherits the fast convergence speed of message passing and, meanwhile, fully leverages the image prior through the integration of score networks. We exploit the fact that, at each iteration, message passing algorithms construct a pseudo AWGN observation of the ground-truth image. We repurpose the score function for denoising the AWGN observation through Tweedie’s formula [12], achieving per-iteration minimum mean squared error (MMSE) performance with a single call of the score network. We further relate the posterior variance calculation to the second-order generalization of Tweedie’s formula, which depends on the Hessian of the log density (i.e., the second-order score function) [13]. We propose matching the second-order score function to obtain the posterior variance, facilitating message updates in a turbo-type message passing algorithm, hence the name score-based turbo message passing (STMP). Experimental results on the FFHQ dataset [14] demonstrate that STMP strikes a much better tradeoff between performance and complexity than conventional message passing, regularized linear regression, and score-based posterior sampling schemes.
Figure 1: A modularized representation of the Turbo-CS framework [15].
II Review of Turbo-CS Framework
The Turbo-CS framework [15], also known as vector AMP (VAMP) [16], is a well-established method for approximating the MMSE estimator $\mathbb{E}[\mathbf{x}|\mathbf{y}]$. This framework can be derived using various approaches such as expectation consistent (EC) approximation and expectation propagation (EP). As illustrated in Fig. 1, Turbo-CS consists of two modules and admits a modularized interpretation that resembles the turbo decoder. In particular, module A is a linear MMSE (LMMSE) estimator of $\mathbf{x}$ exploiting the likelihood $p(\mathbf{y}|\mathbf{x})$ and the messages from module B. Module B is an MMSE denoiser incorporating the prior distribution $p(\mathbf{x})$ and the messages from module A. The two modules are executed iteratively to refine the estimate.
Module A calculates the posterior estimate of $\mathbf{x}$ by combining the likelihood $p(\mathbf{y}|\mathbf{x})=\mathcal{N}(\mathbf{y};\mathbf{A}\mathbf{x},\delta_{0}^{2}\mathbf{I})$ with the pseudo Gaussian prior $\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}}\mathbf{I}\big)$ as
$\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf A}^{\textit{post}},v_{\sf A}^{\textit{post}}\mathbf{I}\big)\propto\mathcal{N}(\mathbf{y};\mathbf{A}\mathbf{x},\delta_{0}^{2}\mathbf{I})\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}}\mathbf{I}\big),$ (2)
where
@@ II Review of Turbo-CS Framework
Eqns. (3) and (4) correspond to the posterior mean and variance from the LMMSE estimator, respectively. The posterior distribution cannot be directly passed to module B due to the correlation issue [17]. Instead, we compute the extrinsic message by excluding the contribution of the prior message as
$\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf A}^{\textit{ext}},v_{\sf A}^{\textit{ext}}\mathbf{I}\big)\propto\frac{\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf A}^{\textit{post}},v_{\sf A}^{\textit{post}}\mathbf{I}\big)}{\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}}\mathbf{I}\big)},$ (5)
where the extrinsic variance and mean are respectively given by
$v_{\sf A}^{\textit{ext}}$ $=\left(\frac{1}{v_{\sf A}^{\textit{post}}}-\frac{1}{v_{\sf A}^{\textit{pri}}}\right)^{-1},$ (6)
$\mathbf{x}_{\sf A}^{\textit{ext}}$ $=v_{\sf A}^{\textit{ext}}\left(\frac{\mathbf{x}_{\sf A}^{\textit{post}}}{v_{\sf A}^{\textit{post}}}-\frac{\mathbf{x}_{\sf A}^{\textit{pri}}}{v_{\sf A}^{\textit{pri}}}\right).$ (7)
The extrinsic message of module A serves as the input to module B, i.e., $\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf B}^{\textit{pri}},v_{\sf B}^{\textit{pri}}\mathbf{I}\big)=\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf A}^{\textit{ext}},v_{\sf A}^{\textit{ext}}\mathbf{I}\big)$. Module B assumes a pseudo likelihood $p\big(\mathbf{x}_{\sf B}^{\textit{pri}}|\mathbf{x}\big)=\mathcal{N}\big(\mathbf{x}_{\sf B}^{\textit{pri}};\mathbf{x},v_{\sf B}^{\textit{pri}}\mathbf{I}\big)$. That is, $\mathbf{x}_{\sf B}^{\textit{pri}}$ is modeled as an AWGN observation of the ground-truth $\mathbf{x}$:
$\mathbf{x}_{\sf B}^{\textit{pri}}=\mathbf{x}+\mathbf{w},~~\mathbf{w}\sim\mathcal{N}\big(\mathbf{0},v_{\sf B}^{\textit{pri}}\mathbf{I}\big).$ (8)
This assumption is justified by the state evolution (SE) analysis in [18], showing that the error vector $\mathbf{x}_{\sf B}^{\textit{pri}}-\mathbf{x}$ is asymptotically i.i.d. Gaussian in the large system limit with right-rotationally invariant $\mathbf{A}$. Module B calculates the posterior estimate of $\mathbf{x}$ by combining the prior $p(\mathbf{x})$ with the pseudo Gaussian likelihood $p\big(\mathbf{x}_{\sf B}^{\textit{pri}}|\mathbf{x}\big)=\mathcal{N}\big(\mathbf{x}_{\sf B}^{\textit{pri}};\mathbf{x},v_{\sf B}^{\textit{pri}}\mathbf{I}\big)$ following the EP rule as
$\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf B}^{\textit{post}},v_{\sf B}^{\textit{post}}\mathbf{I}\big)=\mathrm{proj}\big[p(\mathbf{x})\mathcal{N}\big(\mathbf{x}_{\sf B}^{\textit{pri}};\mathbf{x},v_{\sf B}^{\textit{pri}}\mathbf{I}\big)\big],$ (9)
where $\mathrm{proj}[p(\cdot)]$ denotes the projection of $p(\cdot)$ to a Gaussian density with matched first- and second-order moments. Eqn. (9) can be interpreted as the MMSE denoising under the AWGN observation model (8):
$\mathbf{x}_{\sf B}^{\textit{post}}$ $=\mathbb{E}\big[\mathbf{x}\big|\mathbf{x}_{\sf B}^{\textit{pri}}\big],$ (10)
$v_{\sf B}^{\textit{post}}$ $=\frac{1}{N}\mathrm{tr}\big(\mathrm{Cov}\big[\mathbf{x}\big|\mathbf{x}_{\sf B}^{\textit{pri}}\big]\big).$ (11)
The extrinsic message of module B is calculated similar to that of module A as
$v_{\sf B}^{\textit{ext}}$ $=\left(\frac{1}{v_{\sf B}^{\textit{post}}}-\frac{1}{v_{\sf B}^{\textit{pri}}}\right)^{-1},$ (12)
$\mathbf{x}_{\sf B}^{\textit{ext}}$ $=v_{\sf B}^{\textit{ext}}\left(\frac{\mathbf{x}_{\sf B}^{\textit{post}}}{v_{\sf B}^{\textit{post}}}-\frac{\mathbf{x}_{\sf B}^{\textit{pri}}}{v_{\sf B}^{\textit{pri}}}\right).$ (13)
Finally, the extrinsic message of module B flows leftward to module A, serving as the prior message of module A in the next iteration, i.e., $\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf A}^{\textit{pri}},v_{\sf A}^{\textit{pri}}\mathbf{I}\big)=\mathcal{N}\big(\mathbf{x};\mathbf{x}_{\sf B}^{\textit{ext}},v_{\sf B}^{\textit{ext}}\mathbf{I}\big)$.
@@
The main challenge in implementing the Turbo-CS algorithm arises from the MMSE denoising part in module B, primarily due to i) the lack of a tractable prior distribution and ii) the high-dimensional integrals involved in calculating the posterior mean and variance. To sidestep these difficulties, various off-the-shelf image denoisers (e.g., TV, SURE-LET [4], BM3D [5]) are plugged into message passing algorithms in place of the principled MMSE denoiser. These plug-and-play denoisers typically rely on generic or hand-crafted priors, which could lead to substantial imperfection in compressive imaging applications due to oversimplification and misspecification of the prior distribution. Motivated by the above, we propose a score-based turbo message passing (STMP) algorithm that exploits the score-based generative models as an accurate and expressive image prior, facilitating compressive image recovery from a Bayesian perspective.
III Score-Based MMSE Denoising
Theorem 1 (Tweedie’s formula and its second-order generalization [12])